OpenStudy (anonymous):

Use the substitution method to solve the linear system. 4x+y=9 y=-7

OpenStudy (anonymous):

Can you put in y as -7 into the first equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Substitute y=-7 into the equation; 4x + (-7) = 9 rearrange for x; 4x = 9 + 7 => x = 16/4 => x=4

OpenStudy (amistre64):

4x+y=9 y=-7 4x+y=9 -y=7 -------- 4x = 16